# InstructPipe: Building Visual Programming Pipelines with Human Instructions Using LLMs

### v2

@@ Abstract.
copyright: rightsretainedjournalyear: 2024copyright: rightsretaineddoi: XX.XXXX/XXXXXXXisbn: XXX-X-XXXX-XXXX-X/XX/XXccs: Computing methodologies Visual analyticsccs: Software and its engineering Visual languagesccs: Computing methodologies Machine learning
@@ 1. Introduction
Despite the development of visual programming platforms in various domains, we observed that existing systems share one similar characteristic: users usually initiate a creative process in the workspace “from scratch”. This implies that users need to 1) select nodes, 2) ideate the pipeline structure, and finally, 3) connect nodes within a completely empty workspace. For users unfamiliar with a particular visual programming platform, such processes can be overwhelming, degrading their overall programming experience. Providing pipeline templates may reduce on-boarding efforts (Chen et al., 2024; Fogarty et al., 2001), but the templates inherently lack flexibility and are not easily adaptable to users’ specific needs. Similar issues also arise when users write programs using text-based editors (there exist many built-in functions in a particular programming language and multiple variables in a program), but advances in LLM assistants show that such challenges can be effectively reduced. For example, GitHub Copilot (GitHub, 2023) enables users to generate code by simply describing users’ requirements in natural language. Even though the generated code is not absolutely correct, the AI assistance usually finishes a large portion of the task, and programmers may only need to make a few edits to achieve a correct result (Chris Perry, 2023; Jupyter-ai, 2024). To this end, we raise the following question that motivates our work: How can we build visual programming assistants to accelerate design and prototyping of ML pipelines?
@@
This paper introduces InstructPipe, a visual programming AI assistant that enables ML pipeline generation and design through natural language instructions. InstructPipe facilitates node connection and selection, allowing users to focus on more creative tasks like parameter tuning and interactive analysis within the visual programming workspace. We focus our AI assistant exploration on ML-based pipelines, and therefore implemented InstructPipe as an extension to Visual Blocks (Du et al., 2023a), a visual programming system for prototyping ML pipelines. One major technical challenge in implementing InstructPipe lies in the lack of visual programming data, making it impractical to finetune a dedicated code-LLM similar to how developers build text-editor-based copilots (GitHub, 2023; Chris Perry, 2023; Jupyter-ai, 2024). We addressed this issue by decomposing the generation process into three steps (Fig. 1). InstructPipe’s first LLM module scopes the potentially useful nodes, while the second LLM module generates pseudocode for a pipeline. InstructPipe then parses the pseudocode and renders the pipeline in the node-graph editor to facilitate further user interaction. Our technical evaluation suggests that InstructPipe reduces the necessary user interactions by $81.1\%$ when users select and connect nodes, compared to building them from scratch. This can potentially streamline the development process, and allows users to focus on more novice-friendly interactions like parameter-tuning and human-in-the-loop verification. Our system evaluation with 16 participants demonstrated that InstructPipe significantly reduced users’ workload in their creative process. Qualitative results further reveal that InstructPipe effectively supports novices’ on-boarding experience of visual programming systems and allows them to easily prototype concepts for various purposes. In our experiments, we also observed new challenges caused by human cognitive characteristics, and propose future technical directions towards open-ended AI prototyping assistants.
@@ 2.1. Visual Programming
Advancements in AI has introduced many repositories of advanced ML models (Shen et al., 2024; HuggingFace, 2022), and an increasing number of researchers are exploring AI chains (Wu et al., 2022b; LangChain, 2023). This progress has motivated HCI researchers to design and build a range of visual programming interfaces to support the AI development process (Wu et al., 2022a; LangFlow, 2023; ComfyUI, 2023). For example, ChainForge is a web-based platform for developers to explore various LLM-related configuration and designs in a wide range of applications (Arawjo et al., 2024). Visual Blocks enables creation and interaction of advanced ML pipelines that can leverage state-of-the-art computer vision and computer graphics models in the browser (Du et al., 2023a).
@@ 2.2. Interactive Systems with LLMs
Many HCI researchers integrated LLMs in conventional interactive systems and demonstrated that such enhanced machine intelligence can provide new user experiences (Peng et al., 2023; Park et al., 2022; Liu et al., 2023a; Wang et al., 2023b). This research principle is widely applied in many downstream HCI applications including visualization (Wang et al., 2023a; Shen et al., 2023), explainable AI (Wu et al., 2021; Wang et al., 2024), and social science (Park et al., 2023; Liu et al., 2023b). For example, Chen et al. (Chen et al., 2023) utilized LLMs to bridge low-level sensor information with high-level human requests. Experiments showed that such connection allows users to “construct their personalized contexts [for an intelligent system] more quickly, accurately, and naturally”. To interface human intention with machine operations, researchers typically utilized LLMs by following the ReAct (reasoning and acting) paradigm (Yao et al., 2023). For example, Park et al. (Park et al., 2023) simulated human behaviors in an artifical social system by leveraging LLMs as intelligent agents that can perceiving the environment, planning their behaviors, and acting in the environment. Automated Visualization (AutoViz) researchers employ LLMs for data analysis and reasoning for presenting the visualization (Narechania et al., 2021; Mitra et al., 2022; Sah et al., 2024). For example, LIDA features four modules in the visualization pipeline to 1) summarize a structured dataset, 2) explore user’s goal, 3) generate code for visualization, and 4) render visualization (Dibia, 2023). ChartGPT further constructs a dedicated dataset for chart visualization, and finetunes an LLM for fully automating the data visualization pipeline (Tian et al., 2024).
InstructPipe extends the application of ReAct-like LLM frameworks (Tian et al., 2024; Dibia, 2023) to visual programming and demonstrates its effectiveness to support rapid prototyping with lower user workload. Additionally, introducing visual programming to the ReAct framework showcases an interface solution for human-AI collaboration. That being said, our work values partially correct AI generation, though the previous literature considers it as a complete generation failure (Dibia, 2023; Gupta and Kembhavi, 2023). We render the imperfect results in a visual programming platform, so that even novices can intuitively stand on the shoulder of the AI generation for their work.
@@ 3. InstructPipe
InstructPipe is an AI assistant that enables users to generate a visual programming pipeline by simply providing text-based instructions. We implemented InstructPipe on Visual Blocks (Du et al., 2023a), a visual programming system for prototyping ML pipelines.
@@ 3.1. User Workflow
To generate a pipeline, users first click the “InstructPipe” button in the top-right corner of the interface (2(b)). The system then activates a simple dialog (2(a)) in which users provide a description and a tag for their desired pipeline. The tag can be “language”, “visual”, or “multimodal”, and helps guide the pipeline generation. After users click the “Submit” button, InstructPipe generates a visual pipeline in the node-graph editor. More specifically, InstructPipe generates a directed acyclic graph (DAG) of a visual programming pipeline. This implies that it uses default node parameters (e.g., the “temporature” or “max_tokens” value of an LLM node). Therefore, after the generation, the user needs to 1) finish the graphic structure if necessary, and 2) perform parameter tuning as well as human-in-the-loop evaluation of the pipeline quality interactively in the visual programming platform. As we will show in our evaluation, this new human-AI collaboration approach reduces users’ workload on the technical portion of the visual programming tasks (selecting and connecting nodes), and thus provide a more novice-friendly experiences for technical visual programming platforms.
@@ 3.2. Primitive Nodes
InstructPipe supports 27 primitive nodes in Visual Blocks. We achieved this node library of InstructPipe by filtering out nodes without explicit definition of their functions [^1] Note that Visual Blocks is a system that is actively being updated, and there are more nodes now.. For example, ‘TFLite model runner’ is an implicitly defined node: the user needs to input a tensorflow hub link to define its functionality. As we mentioned previously, InstructPipe focuses on generating a DAG and leaves the parameter tuning task to users. Adding such implicit nodes without clear definition of the functionality can easily confuse our AI assistant in the generation process, and thus we decide to exclude these nodes in the node library of InstructPipe.
@@
We distributed 20 processor nodes based on the data type of its I/O edges and visualized it in Fig. 3. For example. “Google Web Search” takes “Texts” information as input and output new “Texts”, and “OCR” takes an image (vision-based information) as input and output “Texts”. “Features” in Fig. 3 indicates a wide range of intermediate data format used in ML pipelines, including segmentation masks, a pose landmarks, URLs and etc. As shown in the matrix, InstructPipe contains a wide range of nodes that supports the creation of complex ML pipelines. Compared to related work that automates ad hoc ML inferences in specific use scenarios (Gupta and Kembhavi, 2023; Surís et al., 2023), InstructPipe makes one more step towards the open-ended assistants with a more diverse set of primitive nodes. Further extending our node library can effectively empower the capability of our AI assistant, which we leave as critical future work. In the current implementation of InstructPipe, we focus on demonstrating its capability based on our focused node library, and explore what new experiences this AI assistant can bring to our users.
@@ 4. Pipeline Generation from Instructions
InstructPipe leverages LLMs to generate visual programming pipelines from text instructions. There are two prevailing approaches for LLM-customization, fine-tuning (Qin et al., 2023b; Liu et al., 2023a), and few-shot prompting (Park et al., 2023; Gupta and Kembhavi, 2023). Fine-tuning would require a substantial volume of annotated data, with pairs of pipelines and prompts, and it is hard to achieve for a specific visual programming platform. Additionally, a growing list of nodes would consistently require 1) new data annotation and 2) retraining the model, making this approach less sustainable. In comparison, few-shot prompting is a more practical approach for prototyping an interaction concept to understand the new experience it would bring to the community (Gupta and Kembhavi, 2023; Wei et al., 2023; Yao et al., 2023). One major challenge of applying LLMs in visual programming AI assistant lies in designing efficient prompts that fit within a reasonable number of tokens. Even though we focus our exploration on 27 nodes, the node configuration file alone includes 8200 tokens. Further formulating pipeline examples as in-context few-shot examples would result in a combinatorial explosion, causing an overwhelming number of tokens in the prompt.
To this end, we implement InstructPipe with a two-stage LLM refinement prompting strategy, followed by a pseudocode interpretation step to render a pipeline.
@@
Fig. 1 illustrates the high-level workflow of the InstructPipe implementation. InstructPipe leverages two LLM modules (highlighted in red); a Node Selector (section 4.2), and a Code Writer (section 4.3). Given a user instruction and a pipeline tag, we first devise the Node Selector to identify a list of potential nodes that would be used according to the instruction. In the Node Selector, we prompt the LLM with a very brief description of each node, aiming to filter out unrelated nodes for a target pipeline. The selected nodes and the original user input (the prompt and the tag) are then fed into the Code Writer, which generates pseudocode for the desired pipeline. In Code Writer, we provide the LLM with detailed description and examples of each selected nodes to ensure the LLM has extensive context for each candidate node. Finally, we employ a Code Interpreter to parse the pseudocode and render a visual programming pipeline for the user to interact with.
@@ 4.1. Pipeline Representation
The Visual Blocks system takes a JSON-format data as input and renders a directed acyclic graph (DAG) in the visual programming workspace. Therefore, the ultimate goal of InstructPipe is to generate the JSON file; however, directly generating the long JSON file is computationally expensive. For example, the JSON file for rendering the pipeline in 4(a) contains approximately 2.8k tokens. To address this issue, we utilize the pseudocode representation of a DAG, and define this token-efficient data format as the output data format of our LLM module. 4(b) shows the corresponding pseudocode representation of the pipeline in 4(a), and the it only contains 123 tokens. The pseudocode representation simply stores the DAG information of a visual programming pipeline without other information such as node parameters (e.g., the “max_tokens” configuration of an LLM module) and the layouts of the nodes. This indicates that InstructPipe leaves the task of node parameter tuning to the user, which we believe is a more novice-friendly task, and focuses on providing technical assistants on selecting and connecting nodes.
In the following content, we provide detailed explanation on the pseudocode design and implementation. As we mentioned above, Fig. 4 provides an example of a pipeline (4(a)) and its corresponding pseudocode (4(b)). The syntax design is inspired by TypeScript, and the overall structure is inspired by how academic papers present pseudocode (Zhang et al., 2021) in an algorithm block. In 4(b), we highlight the first line under the processor module (i.e., the operation of the PaLI node) in different colors, representing four different components in the programming language. “pali_1” is the unique node ID. The green symbol after the colon, i.e., “pali”, specifies the node type. In this example, node ID “$pali\_1$” is a “pali” node. The arguments in brackets, i.e., “image=input_image_1, prompt=input_text_1”, specifies the input variables (or input edges in the graph) of this node. “pali_1_out” represents the output variable name. For input nodes, the output variable name is the same as the node id, so we do not annotate the output variable with a separate name (e.g., “input_image_1: input_image()” instead of “input_image_1 = input_image_1: input_image()”). Note that InstructPipe generates texts (i.e., the node parameter) in the “input text” node. Therefore, the argument in “text=“caption this image in detail”” does not indicate that the “input_text” node accepts input edges, but accepts the node parameter input as a special case.

### v3

# InstructPipe: Generating Visual Blocks Pipelines with Human Instructions and LLMs

@@ Abstract.
journalyear: 2025copyright: ccconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3713905isbn: 979-8-4007-1394-1/2025/04ccs: Computing methodologies Visual analyticsccs: Software and its engineering Visual languagesccs: Computing methodologies Machine learning
@@ 1. Introduction
Despite the development of visual programming platforms in various domains, we observed that existing systems share one similar characteristic: users usually initiate a creative process in the workspace “from scratch”. This implies that users need to 1) select nodes, 2) ideate the pipeline structure, and finally, 3) connect nodes within a completely empty workspace. As was also highlighted in existing literature in programming tools (Zhang et al., 2024, 2023a), such processes can easily overwhelm users, especially those who are unfamiliar with a particular visual programming platform. Providing pipeline templates may reduce on-boarding efforts (Chen et al., 2024; Fogarty et al., 2001), but the templates inherently lack flexibility and are not easily adaptable to users’ specific needs. Similar issues also arise when users write programs using text-based editors (there exist many built-in functions in a particular programming language and multiple variables in a program), but advances in LLM assistants show that such challenges can be effectively reduced. For example, GitHub Copilot (GitHub, 2023) enables users to generate code by simply describing users’ requirements in natural language. Even though the generated code is not absolutely correct, the AI assistance usually finishes a large portion of the task, and programmers may only need to make a few edits to achieve a correct result (Chris Perry, 2023; Jupyter-ai, 2024). To this end, we raise the following question that motivates our work: How can we build visual programming assistants to accelerate the design and prototyping of ML pipelines?
@@
This paper introduces InstructPipe, a visual programming AI assistant that enables ML pipeline generation and design through natural language instructions. InstructPipe facilitates node connection and selection, allowing users to focus on more creative tasks like parameter tuning and interactive analysis within the visual programming workspace. We focus our AI assistant exploration on ML-based pipelines, and therefore implemented InstructPipe as an extension to Visual Blocks (Du et al., 2023a), a visual programming system for prototyping ML pipelines. One major technical challenge in implementing InstructPipe lies in the lack of visual programming data, making it impractical to finetune a dedicated code-LLM similar to how developers build text-editor-based copilots (GitHub, 2023; Chris Perry, 2023; Jupyter-ai, 2024). We addressed this issue by decomposing the generation process into three steps (Figure 1). InstructPipe’s first LLM module scopes the potentially useful nodes, while the second LLM module generates pseudocode for a pipeline. InstructPipe then parses the pseudocode and renders the pipeline in the node-graph editor to facilitate further user interaction. Our technical evaluation suggests that InstructPipe reduces the necessary user interactions by $81.1\%$ when users select and connect nodes, compared to building them from scratch. This can potentially streamline the development process, and allows users to focus on more novice-friendly interactions like parameter-tuning and human-in-the-loop verification. Our system evaluation with 16 participants demonstrated that InstructPipe significantly reduced users’ workload in their creative process. Qualitative results further reveal that InstructPipe effectively supports novices’ on-boarding experience of visual programming systems and allows them to easily prototype concepts for various purposes. In our experiments, we also observed new challenges caused by human cognitive characteristics, and proposed future technical directions towards open-ended AI prototyping assistants.
@@ 2.1. Visual Programming
Advancements in AI have introduced many repositories of advanced ML models (Shen et al., 2024; HuggingFace, 2022), and an increasing number of researchers are exploring AI chains (Wu et al., 2022b; LangChain, 2023). This progress has motivated HCI researchers to design and build a range of visual programming interfaces to support the AI development process (Wu et al., 2022a; LangFlow, 2023; ComfyUI, 2023). For example, ChainForge is a web-based platform for developers to explore various LLM-related configuration and designs in a wide range of applications (Arawjo et al., 2024). Visual Blocks enables creation and interaction of advanced ML pipelines that can leverage state-of-the-art computer vision and computer graphics models in the browser (Du et al., 2023a).
@@ 2.2. Interactive Systems with LLMs
Many HCI researchers integrated LLMs in conventional interactive systems and demonstrated that such enhanced machine intelligence can provide new user experiences (Peng et al., 2023; Park et al., 2022; Liu et al., 2023a; Wang et al., 2023b; Feng et al., 2024). This research principle is widely applied in many downstream HCI applications, including visualization (Wang et al., 2023a; Shen et al., 2023), explainable AI (Wu et al., 2021; Wang et al., 2024), and social science (Park et al., 2023; Liu et al., 2023b). For example, Chen et al. (Chen et al., 2023) utilized LLMs to bridge low-level sensor information with high-level human requests. Experiments showed that such connection allows users to “construct their personalized contexts [for an intelligent system] more quickly, accurately, and naturally”. To interface human intention with machine operations, researchers typically utilized LLMs by following the ReAct (reasoning and acting) paradigm (Yao et al., 2023). For example, Park et al. (Park et al., 2023) simulated human behaviors in an artificial social system by leveraging LLMs as intelligent agents that perceive the environment, plan their behaviors, and act in the environment. Automated Visualization (AutoViz) researchers employ LLMs for data analysis and reasoning for presenting the visualization (Narechania et al., 2021; Mitra et al., 2022; Sah et al., 2024). For example, LIDA features four modules in the visualization pipeline to 1) summarize a structured dataset, 2) explore the user’s goal, 3) generate code for visualization, and 4) render visualization (Dibia, 2023). ChartGPT further constructs a dedicated dataset for chart visualization, and finetunes an LLM for fully automating the data visualization pipeline (Tian et al., 2024).
InstructPipe extends the application of ReAct-like LLM frameworks (Tian et al., 2024; Dibia, 2023) to visual programming and demonstrates its effectiveness to support rapid prototyping with lower user workload. Additionally, introducing visual programming to the ReAct framework showcases an interface solution for human-AI collaboration. That being said, our work values partially correct AI generation, though the previous literature considers it as a complete generation failure (Dibia, 2023; Gupta and Kembhavi, 2023). We leverage visual programming as a platform to integrate partially complete AI generations with human interactions, enabling even novices to intuitively collaborate with AI in their creative processes.
@@ 3. InstructPipe
InstructPipe is an AI assistant that enables users to generate a visual programming pipeline by simply providing text-based instructions. We implemented InstructPipe on Visual Blocks (Du et al., 2023a; Zhou et al., 2024), a visual programming system for prototyping ML pipelines.
@@ 3.1. User Workflow
To generate a pipeline, users first click the “InstructPipe” button in the top-right corner of the interface (2(b)). The system then activates a simple dialog (2(a)) in which users provide a description and a tag for their desired pipeline. The tag can be “language”, “visual”, or “multimodal”, and helps guide the pipeline generation. After users click the “Submit” button, InstructPipe generates a visual pipeline in the node-graph editor. More specifically, InstructPipe generates a directed acyclic graph (DAG) of a visual programming pipeline. This implies that it uses default node parameters (e.g., the “temperature” or “max_tokens” value of an LLM node). Therefore, after the generation, the user needs to 1) finish the graphic structure if necessary, and 2) perform parameter tuning as well as human-in-the-loop evaluation of the pipeline quality interactively in the visual programming platform. As we will show in our evaluation, this new human-AI collaboration approach reduces users’ workload on the technical portion of the visual programming tasks (selecting and connecting nodes) and thus provides a more novice-friendly experience for technical visual programming platforms.
@@ 3.2. Primitive Nodes
InstructPipe supports 27 primitive nodes in Visual Blocks. We achieved this node library of InstructPipe by filtering out nodes without explicit definition of their functions [^1] Note that Visual Blocks is a system that is actively being updated, and there are more nodes now.. For example, ‘TFLite model runner’ is an implicitly defined node: the user needs to input a tensorflow hub link to define its functionality. As we mentioned previously, InstructPipe focuses on generating a DAG and leaves the parameter-tuning task to users. Adding such implicit nodes without a clear definition of the functionality can easily confuse our AI assistant in the generation process, and thus, we decide to exclude these nodes in the node library of InstructPipe.
@@
We distributed 20 processor nodes based on the data type of its I/O edges and visualized it in Figure 3. For example. “Google Web Search” takes “Texts” information as input and output new “Texts”, and “OCR” takes an image (vision-based information) as input and output “Texts”. “Features” in Figure 3 indicates a wide range of intermediate data formats used in ML pipelines, including segmentation masks, pose landmarks, URLs and etc. As shown in the matrix, InstructPipe contains a wide range of nodes that support the creation of complex ML pipelines. Compared to related work that automates ad hoc ML inferences in specific use scenarios (Gupta and Kembhavi, 2023; Surís et al., 2023), InstructPipe makes one more step towards the open-ended assistants with a more diverse set of primitive nodes. Further extending our node library can effectively empower the capability of our AI assistant, which we leave as critical future work. In the current implementation of InstructPipe, we focus on demonstrating its capability based on our focused node library and explore what new experiences this AI assistant can bring to our users.
@@ 4. Pipeline Generation from Instructions
InstructPipe leverages LLMs to generate visual programming pipelines from text instructions. There are two prevailing approaches for LLM-customization, fine-tuning (Qin et al., 2023b; Liu et al., 2023a), and few-shot prompting (Park et al., 2023; Gupta and Kembhavi, 2023). Fine-tuning would require a substantial volume of annotated data, with pairs of pipelines and prompts, and it is hard to achieve for a specific visual programming platform. Additionally, a growing list of nodes would consistently require 1) new data annotation and 2) retraining the model, making this approach less sustainable. In comparison, few-shot prompting is a more practical approach for prototyping an interaction concept to understand the new experience it would bring to the community (Gupta and Kembhavi, 2023; Wei et al., 2023; Yao et al., 2023). One major challenge of applying LLMs in visual programming AI assistants lies in designing efficient prompts that fit within a reasonable number of tokens. Even though we focus our exploration on 27 nodes, the node configuration file alone includes 8200 tokens. Further formulating pipeline examples as in-context few-shot examples would result in a combinatorial explosion, causing an overwhelming number of tokens in the prompt.
@@
To this end, we implement InstructPipe with a two-stage LLM refinement prompting strategy, followed by a pseudocode interpretation step to render a pipeline. Figure 1 illustrates the high-level workflow of the InstructPipe implementation. InstructPipe leverages two LLM modules (highlighted in red); a Node Selector (section 4.2), and a Code Writer (section 4.3). Given a user instruction and a pipeline tag, we first devise the Node Selector to identify a list of potential nodes that would be used according to the instruction. In the Node Selector, we prompt the LLM with a very brief description of each node, aiming to filter out unrelated nodes for a target pipeline. The selected nodes and the original user input (the prompt and the tag) are then fed into the Code Writer, which generates pseudocode for the desired pipeline. In Code Writer, we provide the LLM with detailed descriptions and examples of each selected node to ensure the LLM has extensive context for each candidate node. Finally, we employ a Code Interpreter to parse the pseudocode and render a visual programming pipeline for the user to interact with.
@@ 4.1. Pipeline Representation
The Visual Blocks system takes JSON-format data as input and renders a directed acyclic graph (DAG) in the visual programming workspace. Therefore, the ultimate goal of InstructPipe is to generate the JSON file; however, directly generating the long JSON file is computationally expensive. For example, the JSON file for rendering the pipeline in 4(a) contains approximately 2.8k tokens. To address this issue, we utilize the pseudocode representation of a DAG, and define this token-efficient data format as the output data format of our LLM module. 4(b) shows the corresponding pseudocode representation of the pipeline in 4(a), and the it only contains 123 tokens. The pseudocode representation simply stores the DAG information of a visual programming pipeline without other information such as node parameters (e.g., the “max_tokens” configuration of an LLM module) and the layouts of the nodes. This indicates that InstructPipe leaves the task of node parameter tuning to the user, which we believe is a more novice-friendly task, and focuses on providing technical assistance on selecting and connecting nodes.
In the following content, we provide detailed explanation on the pseudocode design and implementation. As we mentioned above, Figure 4 provides an example of a pipeline (4(a)) and its corresponding pseudocode (4(b)). The syntax design is inspired by TypeScript, and the overall structure is inspired by how academic papers present pseudocode (Zhang et al., 2021) in an algorithm block. In 4(b), we highlight the first line under the processor module (i.e., the operation of the PaLI node) in different colors, representing four different components in the programming language. “pali_1” is the unique node ID. The green symbol after the colon, i.e., “pali”, specifies the node type. In this example, node ID “$pali\_1$” is a “pali” node. The arguments in brackets, i.e., “image=input_image_1, prompt=input_text_1”, specifies the input variables (or input edges in the graph) of this node. “pali_1_out” represents the output variable name. For input nodes, the output variable name is the same as the node id, so we do not annotate the output variable with a separate name (e.g., “input_image_1: input_image()” instead of “input_image_1 = input_image_1: input_image()”). Note that InstructPipe generates texts (i.e., the node parameter) in the “input text” node. Therefore, the argument in “text=“caption this image in detail”” does not indicate that the “input_text” node accepts input edges, but accepts the node parameter input as a special case.